A family has two cars. The first car has a fuel efficiency of 15 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 2000 miles, for a total gas consumption of 75 gallons. How many gallons were consumed by each of the two cars that week?
Answer by Theo(13342)

first car gets 15 miles per gallon.
second car gets 40 miles per gallon.
both cars went a total of 2000 miles for a total gas consumption of 75 gallons.

miles driven by first car is x.
miles driven by second car is y.

gallons consumed by first car is x / 15.
gallons consumed by second car is y / 40.

you have 2 equations.

x + y = 2000
x/15 + y/40 = 75

use the first equation to solve for y.

you get y = 2000 - x

replace y with 2000 - x in the second equation to get:

x/15 + (2000 - x) / 40 = 75

remove the fractions by multiplying both sides of the equation by 120.

you get:

120 * x / 15 + 120 * (2000 - x) / 40 = 75 * 120

simplify to get:

8 * x + 3 * (2000 - x) = 9000

simplify further by removing parentheses to get:

8 * x + 6000 - 3 * x = 9000

subtract 6000 from both sides of the equation to get:

8 * x - 3 * x = 3000

combine like terms to get:

5 * x = 3000

divide both sides of the equation by 5 to get:

x = 600

since y = 2000 - x, this means that:

y = 1400

the first car traveled 600 miles.
the second car traveled 1400 miles.

600 + 1400 = 2000
600 / 15 + 1400 / 40 = 40 + 35 = 75

numbers check out.

your solution is the the first car traveled 600 miles at 15 miles per gallon and the second car traveled 1400 miles at 40 miles per gallon.